Assume the absolute zero apparatus is initially open to the air. The volume of a sphere is: V = 4πr3/3

a) Assume that air pressure is Po = 1.00 atm and the air temperature is To = 22.8o C. How many moles of air will the sphere hold?

n air = ___ moles

b) Suppose now that the apparatus, initially open, is now sealed with the air inside. If the apparatus is now heated to to 665o C, find the pressure inside the apparatus now.

P = ___ atm

c) Assume now that, while maintaining a temperature of 665o, the sphere is injected with another 0.029 moles of air. What is the new pressure inside the ball?

P = ___ atm
Answer #1

a)

we know that

PV = nRT

n = PV/RT

n = 4πr3/3 (1/0.082x295.8

b)

here v=cosnrant

so

P2/P1 = T2/T1

P2 = 3.17atm

c)

P2/P1 = T2 x n2/T1 n1